OpenStudy (anonymous):

When you are given the original graph of F(x) and they ask you to get the graph f(x+3) do I just move all teh points to the right 3 times?

OpenStudy (anonymous):

no, it depends on the function......

OpenStudy (anonymous):

say f(x) = x^2 then for x = 1, f(x) = 1^2 = 1 but f(x+3) = (x+3)^2 then for x = 1 f(x+3) = (1+3)^2 = 16

jimthompson5910 (jim_thompson5910):

f(x+3) shifts everything to the left (not right) 3 units